OpenStudy (anonymous):

|x − 0.525| = 0.01Show your work in solving the equation to determine the minimum and maximum diameters of the drill bits. Need answer asap plz :)

OpenStudy (aum):

If |A| = B it means A = B or A = -B If |x − 0.525| = 0.01 it means x − 0.525 = 0.01 or x − 0.525 = -0.01 Solve for x in both cases. The lower value will be the minimum. The higher value will be the maximum.

OpenStudy (anonymous):

Thanks. :)